OpenStudy (anonymous):

-2^4+50^0+(6)^-3 -16+1-216 -231 is this right?

Parth (parthkohli):

\(-2^4 = -2 \times -2\times -2 \times -2 = 4 \times 4 = 16 \) Hmm...the second step

Parth (parthkohli):

\( \color{Black}{\Rightarrow -2^4 \ne -16 }\)

OpenStudy (anonymous):

thanks. what about the -216 is that right?

Parth (parthkohli):

\( \color{Black}{\Rightarrow \large 6^{-3} = {1 \over 6^3} = {1 \over 216} }\)

OpenStudy (anonymous):

so then it is +216?

OpenStudy (anonymous):

that would make the final answer 233?